### v2

# pySLAM: An Open-Source, Modular, and Extensible Framework for SLAM

Abstract
pySLAM is an open-source Python framework for Visual SLAM, supporting monocular, stereo, and RGB-D cameras. It provides a flexible interface for integrating both classical and modern local features, making it adaptable to various SLAM tasks. The framework includes different loop closure methods, a volumetric reconstruction pipeline, and support for depth prediction models. Additionally, it offers a suite of tools for visual odometry and SLAM applications. Designed for both beginners and experienced researchers, pySLAM encourages community contributions, fostering collaborative development in the field of Visual SLAM.
Introduction
The goal of this document is to present the pySLAM framework, its main features, and usage. pySLAM is a python implementation of a Visual SLAM pipeline that supports monocular, stereo and RGBD cameras. It provides the following features:
  • A wide range of classical and modern local features with a convenient interface for their integration.
  • Various loop closing methods, including descriptor aggregators such as visual Bag of Words (BoW, iBow), Vector of Locally Aggregated Descriptors (VLAD), and modern global descriptors (image-wise descriptors).
  • A volumetric reconstruction pipeline that processes available depth and color images with volumetric integration and provides an output dense reconstruction. This can use TSDF with voxel hashing or incremental Gaussian Splatting.
  • Integration of depth prediction models within the SLAM pipeline. TThese include DepthPro, DepthAnythingV2, RAFT-Stereo, CREStereo, MASt3R, MVDUSt3R, etc.
  • A collection of other useful tools for VO and SLAM.
A convenient entry-point are the following main scripts:
  • main_vo.py combines the simplest VO ingredients without performing any image point triangulation or windowed bundle adjustment. At each step $k$, main_vo.py estimates the current camera pose $C_{k}$ with respect to the previous one $C_{k-1}$. The inter-frame pose estimation returns $[R_{k-1,k},t_{k-1,k}]$ with $\|t_{k-1,k}\|=1$. With this very basic approach, you need to use a ground truth in order to recover a correct inter-frame scale $s$ and estimate a valid trajectory by composing $C_{k}=C_{k-1}[R_{k-1,k},st_{k-1,k}]$. This script is a first start to understand the basics of inter-frame feature tracking and camera pose estimation.
  • main_slam.py adds feature tracking along multiple frames, point triangulation, keyframe management, bundle adjustment, loop closing, dense mapping and depth inference in order to estimate the camera trajectory and build both a sparse and dense map. It’s a full SLAM pipeline and includes all the basic and advanced blocks which are necessary to develop a real visual SLAM pipeline.
  • main_feature_matching.py shows how to use the basic feature tracker capabilities (feature detector + feature descriptor + feature matcher) and allows to test the different available local features.
  • main_depth_prediction.py shows how to use the available depth inference models to get depth estimations from input color images.
  • main_map_viewer.py reloads a saved map and visualizes it. Further details on how to save a map here.
  • main_map_dense_reconstruction.py reloads a saved map and uses a configured volumetric integrator to obtain a dense reconstruction (see here).
pySLAM can be used as flexible baseline framework to experiment with VO/SLAM techniques, local features, descriptor aggregators, global descriptors, volumetric integration and depth prediction. It allows to quickly explore, prototype and develop VO/SLAM pipelines. Users should note that pySLAM is a research framework and a work in progress. It is not optimized for real-time performances.
Enjoy it!
System overview
This section presents some diagram sketches that provide an overview of the main workflow, system components, and class relationships/dependencies. To make the diagrams more readable, some minor components and arrows have been omitted.
SLAM Workflow
SLAM Components
Note: In some case, I used Processes instead of Threads because in Python 3.8 (used by pySLAM) the Global Interpreter Lock (GIL) allows only one thread can execute at a time within a single process. Multiprocessing avoids this limitation and enables better parallelism, though it involves data duplication via pickling. See this related nice post.
Main System Components
Feature Tracker
Feature Matcher
Loop Detector
Depth Estimator
Volumetric Integrator
Usage
Once you have run the script install_all_venv.sh / install_all_conda.sh (follow the instructions above according to your OS), you can open a new terminal and start testing the basic Visual Odometry (VO):
. pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
./main_vo.py
This will process a default KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings), and its groundtruth (available in the same data/videos folder). If matplotlib windows are used, you can stop main_vo.py by focusing/clicking on one of them and pressing the key ‘Q’. As explained above, this very basic script main_vo.py strictly requires a ground truth. Now, with RGBD datasets, you can also test the RGBD odometry with the classes VisualOdometryRgbd or VisualOdometryRgbdTensor (ground truth is not required here).
Similarly, you can test the full SLAM by running main_slam.py:
. pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
./main_slam.py
This will process the same default KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings). You can stop it by focusing/clicking on one of the opened windows and pressing the key ‘Q’ or closing the 3D pangolin GUI.
With both scripts, in order to process a different dataset, you need to update the file config.yaml: * Select your dataset type in the section DATASET (further details in the section Datasets below for further details). This identifies a corresponding dataset section (e.g. KITTI_DATASET, TUM_DATASET, etc). * Select the sensor_type (mono, stereo, rgbd) in the chosen dataset section.
* Select the camera settings file in the dataset section (further details in the section Camera Settings below). * The groudtruth_file accordingly (further details in the section Datasets below and check the files io/ground_truth.py and io/convert_groundtruth.py).
Feature tracking
If you just want to test the basic feature tracking capabilities (feature detector + feature descriptor + feature matcher) and get a taste of the different available local features, run
. pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
\$./main_feature_matching.py
In any of the above scripts, you can choose any detector/descriptor among ORB, SIFT, SURF, BRISK, AKAZE, SuperPoint, etc. (see the section Supported Local Features below for further information).
Some basic examples are available in the subfolder test/loopclosing. In particular, as for feature detection/description, you may want to take a look at test/cv/test_feature_manager.py too.
Loop closing
Different loop closing methods are available, combining aggregation methods and global descriptors.
While running full SLAM, loop closing is enabled by default and can be disabled by setting kUseLoopClosing=False in config_parameters.py. Configuration options can be found in loop_closing/loop_detector_configs.py.
Examples: Start with the examples in test/loopclosing, such as test/loopclosing/test_loop_detector.py.
Vocabulary management
DBoW2, DBoW3, and VLAD require pre-trained vocabularies. ORB-based vocabularies are automatically downloaded in the data folder (see loop_closing/loop_detector_configs.py).
To create a new vocabulary, follow these steps:
  1. 1.
    Generate an array of descriptors: Use the script
    test/loopclosing/test_gen_des_array_from_imgs.py to generate the array of descriptors that will be used to train the new vocabulary. Select your desired descriptor type via the tracker configuration.
  2. 2.
    DBOW vocabulary generation: Train your target DBOW vocabulary by using the script test/loopclosing/test_gen_dbow_voc_from_des_array.py.
  3. 3.
    VLAD vocabulary generation: Train your target VLAD “vocabulary” by using the script test/loopclosing/test_gen_vlad_voc_from_des_array.py.
Vocabulary-free loop closing
Most methods do not require pre-trained vocabularies. Specifically: - iBoW and OBindex2: These methods incrementally build bags of binary words and, if needed, convert (front-end) non-binary descriptors into binary ones. - Others: Methods like HDC_DELF, SAD, AlexNet, NetVLAD, CosPlace, and EigenPlaces directly extract global descriptors and process them using dedicated aggregators, independently from the used front-end descriptors.
As mentioned above, only DBoW2, DBoW3, and VLAD require pre-trained vocabularies.
Double-check your loop detection configuration and verify vocabulary compability
When selecting a loop detection method based on a pre-trained vocabulary(such as DBoW2, DBoW3, and VLAD), ensure the following:
  1. 1.
    The back-end and the front-end are using the same descriptor type (this is also automatically checked for consistency).
  2. 2.
    A corresponding pre-trained vocubulary is available.
For more details, refer to the vocabulary management section.
If you lack a compatible vocabulary for the selected front-end descriptor type, no loop closing method will be used. You have the following options:
  1. a.
    Create and load the vocabulary (refer to the vocabulary management section).
  2. b.
    Choose an *_INDEPENDENT loop detector method, which works with an independent local_feature_manager.
  3. c.
    Select a vocabular-free loop closing method.
See the file loopclosing/loop_detector_configs.py for further details.
Volumetric reconstruction
Dense reconstruction while running SLAM
The SLAM back-end hosts a volumetric reconstruction pipeline. This is disabled by default. You can enable it by setting kUseVolumetricIntegration=True and selecting your preferred method kVolumetricIntegrationType in config_parameters.py. At present, two methods are available: TSDF and GAUSSIAN_SPLATTING (see dense/volumetric_integrator_factory.py). Note that you need CUDA in order to run GAUSSIAN_SPLATTING method.
At present, the volumetric reconstruction pipeline works with: - RGBD datasets - When a depth estimator is used in the back-end or front-end and a depth prediction/estimation gets available for each processed keyframe.
If you want a mesh as output then set kVolumetricIntegrationExtractMesh=True in config_parameters.py.
Reload a saved sparse map and perform dense reconstruction
Use the script main_map_dense_reconstruction.py to reload a saved sparse map and to perform dense reconstruction by using its posed keyframes as input. You can select your preferred dense reconstruction method directly in the script.
  • To check what the volumetric integrator is doing, run in another shell tail -f logs/volumetric_integrator.log (from repository root folder).
  • To save the obtained dense and sparse maps, press the Save button on the GUI.
Reload and check your dense reconstruction
You can check the output pointcloud/mesh by using CloudCompare.
In the case of a saved Gaussian splatting model, you can visualize it by:
  1. 1.
    Using the superslat editor (drag and drop the saved Gaussian splatting .ply pointcloud in the editor interface).
  2. 2.
    Getting into the folder test/gaussian_splatting and running:
    \$ python test_gsm.py --load <gs_checkpoint_path>
Controlling the spatial distribution of keyframe FOV centers
If you are targeting volumetric reconstruction while running SLAM, you can enable a keyframe generation policy designed to manage the spatial distribution of keyframe field-of-view (FOV) centers. The FOV center of a camera is defined as the backprojection of its image center, calculated using the median depth of the frame. With this policy, a new keyframe is generated only if its FOV center is farther than a predefined distance from the nearest existing keyframe’s FOV center. You can enable this policy by setting the following parameters in the yaml setting:
KeyFrame.useFovCentersBasedGeneration: 1  # compute 3D fov centers of camera frames by using median depth and use their distances to control keyframe generation
KeyFrame.maxFovCentersDistance: 0.2       # max distance between fov centers in order to generate a keyframe
Depth prediction
The available depth prediction models can be utilized both in the SLAM back-end and front-end. - Back-end: Depth prediction can be enabled in the volumetric reconstruction pipeline by setting the parameter kVolumetricIntegrationUseDepthEstimator=True and selecting your preferred kVolumetricIntegrationDepthEstimatorType in config_parameters.py. - Front-end: Depth prediction can be enabled in the front-end by setting the parameter kUseDepthEstimatorInFrontEnd in config_parameters.py. This feature estimates depth images from input color images to emulate a RGBD camera. Please, note this functionality is still experimental at present time.
Refer to the file depth_estimation/depth_estimator_factory.py for further details. Both stereo and monocular prediction approaches are supported. You can test depth prediction/estimation by using the script main_depth_prediction.py.
Notes: * In the case of a monocular SLAM configuration, do NOT use depth prediction in the back-end volumetric integration: The SLAM (fake) scale will conflict with the absolute metric scale of depth predictions. With monocular datasets, enable depth prediction to run in the front-end. - The depth inference may be very slow (for instance, with DepthPro it takes ~1s per image on my machine). Therefore, the resulting volumetric reconstruction pipeline may be very slow.
Saving and reloading
Save the a map
When you run the script main_slam.py (main_map_dense_reconstruction.py): - You can save the current map state by pressing the button Save on the GUI. This saves the current map along with front-end, and backend configurations into the default folder results/slam_state (results/slam_state_dense_reconstruction). - To change the default saving path, open config.yaml and update target folder_path in the section: bash   SYSTEM_STATE:     folder_path: results/slam_state   # default folder path (relative to repository root) where the system state is saved or reloaded
Reload a saved map and relocalize in it
  • A saved map can be loaded and visualized in the GUI by running:
    . pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
    ./main_map_viewer.py  #  Use the --path options to change the input path
  • To enable map reloading and relocalization when running main_slam.py, open config.yaml and set
    SYSTEM_STATE:
      load_state: True                  # flag to enable SLAM state reloading (map state + loop closing state)
      folder_path: results/slam_state   # default folder path (relative to repository root) where the system state is saved or reloaded
Note that pressing the Save button saves the current map, front-end, and backend configurations. Reloading a saved map overwrites the current system configurations to ensure descriptor compatibility.
Trajectory saving
Estimated trajectories can be saved in three different formats: TUM (The Open Mapping format), KITTI (KITTI Odometry format), and EuRoC (EuRoC MAV format). pySLAM saves two types of trajectory estimates:
  • Online: In online trajectories, each pose estimate depends only on past poses. A pose estimate is saved at the end of each front-end iteration on current frame.
  • Final: In final trajectories, each pose estimate depends on both past and future poses. A pose estimate is refined multiple times by LBA windows that cover it and by GBA during loop closures.
To enable trajectory saving, open config.yaml and search for the SAVE_TRAJECTORY: set save_trajectory: True, select your format_type (tum, kitti, euroc), and the output filename. For instance for a tum format output:
SAVE_TRAJECTORY:
  save_trajectory: True
  format_type: kitti             # supported formats: `tum`, `kitti`, `euroc`
  output_folder: results/metrics # relative to pyslam root folder
  basename: trajectory           # basename of the trajectory saving output
SLAM GUI
Some quick information about the non-trivial GUI buttons of main_slam.py:
  • Step: Enter the Step by step mode. Press the button Step a first time to pause. Then, press it again to make the pipeline process a single new frame.
  • Save: Save the map into the file map.json. You can visualize it back by using the script /main_map_viewer.py (as explained above).
  • Reset: Reset SLAM system.
  • Draw Ground Truth: If a ground truth dataset is loaded (e.g., from KITTI, TUM, EUROC, or REPLICA), you can visualize it by pressing this button. The ground truth trajectory will be displayed in 3D and progressively aligned (approximately every 30 frames) with the estimated trajectory. The alignment improves as more samples are added to the estimated trajectory. After ~20 frames, if the button is pressed, a window will appear showing the Cartesian alignment errors (ground truth vs. estimated trajectory) along the axes.
Monitor the logs for tracking, local mapping, and loop closing simultaneously
The logs generated by the modules local_mapping.py, loop_closing.py, loop_detecting_process.py, and global_bundle_adjustments.py are collected in the files local_mapping.log, loop_closing.log, loop_detecting.log, and gba.log, which are all stored in the folder logs. For debugging, you can monitor a parallel flow by running the following command in a separate shell:
\$ tail -f logs/<log file name>
Otherwise, to check all parallel logs with tmux, run:
\$ ./scripts/launch_tmux_logs.sh
To launch slam and check all logs in a single tmux, run:
\$ ./scripts/launch_tmux_slam.sh
Press CTRL+A and then CTRL+Q to exit from tmux environment.
Supported components and models
Supported local features
At present time, the following feature detectors are supported:
The following feature descriptors are supported:
For more information, refer to local_features/feature_types.py file. Some of the local features consist of a joint detector-descriptor. You can start playing with the supported local features by taking a look at test/cv/test_feature_manager.py and main_feature_matching.py.
In both the scripts main_vo.py and main_slam.py, you can create your preferred detector-descritor configuration and feed it to the function feature_tracker_factory(). Some ready-to-use configurations are already available in the file local_features/feature_tracker.configs.py
The function feature_tracker_factory() can be found in the file local_features/feature_tracker.py. Take a look at the file local_features/feature_manager.py for further details.
N.B.: You just need a single python environment to be able to work with all the supported local features!
Supported matchers
See the file local_features/feature_matcher.py for further details.
Supported global descriptors and local descriptor aggregation methods
Local descriptor aggregation methods
NOTE: iBoW and OBIndex2 incrementally build a binary image index and do not need a prebuilt vocabulary. In the implemented classes, when needed, the input non-binary local descriptors are transparently transformed into binary descriptors.
Global descriptors
Also referred to as holistic descriptors:
Different loop closing methods are available. These combines the above aggregation methods and global descriptors. See the file loop_closing/loop_detector_configs.py for further details.
Supported depth prediction models
Both monocular and stereo depth prediction models are available. SGBM algorithm has been included as a classic reference approach.
Supported volumetric mapping methods
  • TSDF with voxel block grid (parallel spatial hashing)
  • Incremental 3D Gaussian Splatting. See here and MonoGS for a description of its backend [18].
Camera Settings
The folder settings contains the camera settings files which can be used for testing the code. These are the same used in the framework ORB-SLAM2 [36]. You can easily modify one of those files for creating your own new calibration file (for your new datasets).
In order to calibrate your camera, you can use the scripts in the folder calibration. In particular: 1. Use the script grab_chessboard_images.py to collect a sequence of images where the chessboard can be detected (set the chessboard size therein, you can use the calibration pattern calib_pattern.pdf in the same folder) 2. Use the script calibrate.py to process the collected images and compute the calibration parameters (set the chessboard size therein)
For more information on the calibration process, see this tutorial [30] or this other link [42].
If you want to use your camera, you have to:
  • Calibrate it and configure WEBCAM.yaml accordingly.
  • Record a video (for instance, by using save_video.py in the folder calibration).
  • Configure the VIDEO_DATASET section of config.yaml in order to point to your recorded video.
Comparison pySLAM vs ORB-SLAM3
For a comparative evaluation, online trajectory estimated by pySLAM vs final trajectories estimated by ORB-SLAM3, see this nice notebook. Note that pySLAM is able to save both online and final pose estimates. On the other end, ORB-SLAM3 pose estimates are saved at the end of the full dataset playback. For further details about online/final trajectories and trajectory saving, see the section Trajectory saving.
Credits
The following is a list of frameworks that inspired or has been integrated into pySLAM. Many thanks to their Authors for their great work.