### v3

# pySLAM: An Open-Source, Modular, and Extensible Framework for SLAM

Abstract
pySLAM is an open-source Python framework for Visual SLAM that supports monocular, stereo, and RGB-D camera inputs. It offers a flexible and modular interface, integrating a broad range of both classical and learning-based local features. The framework includes multiple loop closure strategies, a volumetric reconstruction pipeline, and support for depth prediction models. It also offers a comprehensive set of tools for experimenting with and evaluating visual odometry and SLAM modules. Designed for both beginners and experienced researchers, pySLAM emphasizes rapid prototyping, extensibility, and reproducibility across diverse datasets. Its modular architecture facilitates the integration of custom components and encourages research that bridges traditional and deep learning-based approaches. Community contributions are welcome, fostering collaborative development and innovation in the field of Visual SLAM. This document [^1] You may find an updated version of this document at:
github.com/luigifreda/pyslam/blob/master/docs/tex/document.pdf
presents the pySLAM framework, outlining its main components, features, and usage.
1 Introduction
pySLAM is an open-source framework designed to accelerate research and development in Visual SLAM (Simultaneous Localization and Mapping) and Spatial AI. It originated from a practical need identified during the development of a SLAM/Spatial AI course: The absence of a unified, accessible, and extensible SLAM ecosystem in C++ and Python. The project was driven by several foundational motivations, including the desire to rapidly prototype and evaluate both classical and learning-based features within complete VO/SLAM pipelines, ensure reproducibility across a wide range of datasets, and significantly lower the entry barrier for newcomers to the field.
A key decision behind pySLAM development was the choice of Python as its primary language. Compared to C++, Python enables faster experimentation and benefits from a mature and expanding ecosystem of deep learning tools and research-oriented APIs. This choice aligns with current trends in computer vision and robotics research, where rapid prototyping and flexible experimentation are essential.
However, designing a modern SLAM framework in Python presents several challenges. The SLAM landscape is highly fragmented – across feature types, datasets, programming languages (Python vs. C++), and development environments (virtual environments, Docker, versioning issues). Deep learning integration further complicates matters due to backend fragmentation (e.g., PyTorch vs. TensorFlow). These inconsistencies often lead to reproducibility bottlenecks and high barriers to experimental research, where setting up a functional baseline can become more laborious than the research itself.
To address these issues, pySLAM is built on several key design principles. Modularity is central: components such as local/global features, loop closure mechanisms, semantic mapping, and volumetric integration can be easily swapped or extended. This architecture supports the use of interchangeable baselines, enabling researchers to plug-and-play with different configurations depending on their goals. Extensibility is facilitated through a base-class architecture that allows new components to be integrated with minimal friction, making pySLAM a flexible platform for benchmarking and algorithm development.
In terms of usability, pySLAM emphasizes simplicity in the build and deployment process.. It supports a unified Python environment compatible with conda, pyenv, and pixi, and offers a one-command installation process that covers setup and dataset acquisition. Cross-platform compatibility (Linux, macOS, Windows with WSL2) and full source accessibility further enhance its utility for both academic and industrial use.
Another cornerstone of the framework is standardized dataset handling. By providing a unified interface to multiple datasets, pySLAM encourages reproducibility and reusability. It supports consistent data loading, map state saving/loading, and automatic generation of evaluation metrics – including final and online trajectory estimation – thus enabling robust comparison between algorithms.
Functionally, pySLAM supports the full visual SLAM pipeline with monocular, stereo, and RGB-D configurations. It accommodates both traditional local features and deep-learning-based methods, and includes loop closure detection integrated with optimization backends like g2o and GTSAM. The framework also supports 3D reconstruction via TSDF and emerging techniques like Gaussian Splatting, and integrates deep learning models for depth prediction and semantic scene understanding.
Together, these capabilities position pySLAM as a powerful research toolkit – aiming to be a “Swiss army knife“ for SLAM development, evaluation, and education. By streamlining reproducibility, supporting modular experimentation, and lowering the technical threshold for entry, pySLAM addresses some of the most persistent challenges in contemporary SLAM research.
The objective of this document is to present the pySLAM framework, its main features, and usage [^2] You may find an updated version of this document at:
github.com/luigifreda/pyslam/blob/master/docs/tex/document.pdf
2 Main Features
pySLAM provides a python implementation of a Visual SLAM pipeline that supports monocular, stereo and RGBD cameras. It offers the following features in a single python environment:
  • A wide range of classical and modern local features with a convenient interface for their integration.
  • Various loop closing methods, including descriptor aggregators such as visual Bag of Words (BoW, iBow), Vector of Locally Aggregated Descriptors (VLAD), and modern global descriptors (image-wise descriptors).
  • A volumetric reconstruction pipeline that processes available depth and color images with volumetric integration and provides an output dense reconstruction. This can use TSDF with voxel hashing or incremental Gaussian Splatting.
  • Integration of depth prediction models within the SLAM pipeline. These include DepthPro, DepthAnythingV2, RAFT-Stereo, CREStereo, MASt3R, MVDUSt3R, etc.
  • A suite of segmentation models for semantic understanding of the scene, such as DeepLabv3, Segformer, and dense CLIP.
  • Additional tools for VO (Visual Odometry) and SLAM, with built-in support for both g2o and GTSAM, along with custom Python bindings for features not available in the original libraries
  • Built-in support for over 10 dataset types. The list is currently growing.
pySLAM serves as a flexible baseline framework to experiment with VO/SLAM techniques, local features, descriptor aggregators, global descriptors, volumetric integration, depth prediction and semantic mapping. It allows to explore, prototype and develop VO/SLAM pipelines. pySLAM is a research framework and a work in progress. It is not optimized for real-time performance.
3 Overview
3.1 Main Scritps
A convenient entry point is the following collection of main scripts:
  • main_vo.py combines the simplest VO ingredients without performing any image point triangulation or windowed bundle adjustment. At each step $k$, main_vo.py estimates the current camera pose $C_{k}$ with respect to the previous one $C_{k-1}$. The inter-frame pose estimation returns $[R_{k-1,k},t_{k-1,k}]$ with $\|t_{k-1,k}\|=1$. With this very basic approach, you need to use a ground truth in order to recover a correct inter-frame scale $s$ and estimate a valid trajectory by composing $C_{k}=C_{k-1}[R_{k-1,k},st_{k-1,k}]$. This script is a first start to understand the basics of inter-frame feature tracking and camera pose estimation.
  • main_slam.py adds feature tracking along multiple frames, point triangulation, keyframe management, bundle adjustment, loop closing, dense mapping and depth inference in order to estimate the camera trajectory and build both a sparse and dense map. It’s a full SLAM pipeline and includes all the basic and advanced blocks which are necessary to develop a real visual SLAM pipeline.
  • main_feature_matching.py shows how to use the basic feature tracker capabilities (feature detector + feature descriptor + feature matcher) and allows to test the different available local features.
  • main_depth_prediction.py shows how to use the available depth inference models to get depth estimations from input color images.
  • main_map_viewer.py reloads a saved map and visualizes it. Further details on how to save a map here.
  • main_map_dense_reconstruction.py reloads a saved map and uses a configured volumetric integrator to obtain a dense reconstruction (see here).
  • main_slam_evaluation.py enables automated SLAM evaluation by executing main_slam.py across a collection of datasets and configuration presets (see here).
Figure 1: SLAM workflow.
Figure 2: SLAM components.
3.2 System overview
This section presents some diagram sketches that provide an overview of the main workflow, system components, and class relationships/dependencies. To make the diagrams more readable, some minor components and arrows have been omitted.
SLAM Workflow and Components
Fig. 1 illustrates the SLAM workflow, which is composed of six main parallel processing modules:
  • Tracking: estimates the camera pose for each incoming frame by extracting and matching local features to the local map, followed by minimizing the reprojection error through motion-only Bundle Adjustment (BA). It includes components such as pose prediction (or relocalization), feature tracking, local map tracking, and keyframe decision-making.
  • Local Mapping: updates and refines the local map by processing new keyframes. This involves culling redundant map points, creating new points via temporal triangulation, fusing nearby map points, performing Local BA, and pruning redundant local keyframes.
  • Loop Closing: detects and validates loop closures to correct drift accumulated over time. Upon loop detection, it performs loop group consistency checks and geometric verification, applies corrections, and then launches Pose Graph Optimization (PGO) followed by a full Global Bundle Adjustment (GBA). Loop detection itself is delegated to a parallel process, the Loop Detector, which operates independently for better responsiveness and concurrency.
  • Global Bundle Adjustment: triggered by the Loop Closing module after PGO, this step globally optimizes the trajectory and the sparse structure of the map to ensure consistency across the entire sequence.
  • Volumetric Integration: uses the keyframes, with their estimated poses and back-projected point clouds, to reconstruct a dense 3D map of the environment. This module optionally integrates predicted depth maps and maintains a volumetric representation such as a TSDF [15] or incremental Gaussian Splatting-based volume [37, 21].
  • Semantic Mapping: enriches the SLAM map with dense semantic information by applying pixel-wise segmentation to selected keyframes. Semantic predictions are fused across views to assign semantic labels or descriptors to keyframes and map points. The module operates in parallel, consuming keyframes and associated image data from a queue, applying a configured semantic segmentation model, and updating the map with fused semantic features. This enables advanced downstream tasks such as semantic navigation, scene understanding, and category-level mapping.
The first four modules follow the established PTAM [22] and ORB-SLAM [43] paradigm. Here, the Tracking module serves as the front-end, while the remaining modules operate as part of the back-end.
In parallel, the system constructs two types of maps:
  • a sparse map ${\cal M}_{s}=({\cal K},{\cal P})$, composed of a set of keyframes $\cal K$ and 3D points ${\cal P}$ derived from matched features;
  • a volumetric map (or dense map) ${\cal M}_{v}$, constructed by the Volumetric Integration module, which fuses back-projected point clouds from the keyframes $\cal K$ into a dense 3D model.
To ensure consistency between the sparse and volumetric representations, the volumetric map is updated or re-integrated whenever global pose adjustments occur (e.g., after loop closures).
Fig. 2 details the internal components and interactions of the above modules. In certain cases, processes are employed instead of threads. This is due to Python’s Global Interpreter Lock (GIL), which prevents concurrent execution of multiple threads in a single process. The use of multiprocessing circumvents this limitation, enabling true parallelism at the cost of some inter-process communication overhead (e.g., via pickling). For an insightful discussion, see this related post.
3.3 Main System Components
Figure 3: Feature tracker.
Feature Tracker
The Feature Tracker consists of the following key sub-components:
  • Feature Manager: Manages local feature operations. It includes the FeatureDetector, FeatureDescriptor, and adaptors for pyramid management and image tiling.
    • Feature Detector: Identifies salient and repeatable keypoints in the image, such as corners or blobs, which are likely to be robust under viewpoint and illumination changes.
    • Feature Descriptor: Computes a distinctive descriptor for each detected keypoint, encoding its local appearance to enable robust matching across frames. Examples include ORB [54], SIFT [31], or SuperPoint [14] descriptors.
  • Feature Matcher: Establishes correspondences between features in successive frames (or stereo pairs) by comparing their descriptors or directly inferring matches from image content. Matching can be performed using brute-force, k-NN with ratio test, or learned matching strategies. Refere to Sect. 3.3 for futher details.
Sect. 5.1 lists the supported local feature extractors and detectors.
The diagram in Fig. 3 presents the architecture of the Feature Tracker system. It is structured around a feature_tracker_factory, which instantiates specific tracker types such as LK, DES_BF, DES_FLANN, XFEAT, LIGHTGLUE, and LOFTR. Each tracker type creates a corresponding implementation (e.g., LKFeatureTracker, DescriptorFeatureTracker, etc.), all of which inherit from a common FeatureTracker interface.
The FeatureTracker class is composed of several key sub-components, including a FeatureManager, FeatureDetector, FeatureDescriptor, PyramidAdaptor, BlockAdaptor, and FeatureMatcher. The FeatureManager also encapsulates instances of the detector, descriptor, and adaptors, highlighting the modular and reusable design of the tracking pipeline.
Figure 4: Feature matcher.
Figure 5: Loop detector.
Feature Matcher
The diagram in Fig. 4 illustrates the architecture of the Feature Matcher module. At its core is the feature_matcher_factory, which instantiates matchers based on a specified matcher_type, such as BF, FLANN, XFEAT, LIGHTGLUE, and LOFTR. Each of these creates a corresponding matcher implementation (e.g., BfFeatureMatcher, FlannFeatureMatcher, etc.), all inheriting from a common FeatureMatcher interface.
The FeatureMatcher class encapsulates several configuration parameters and components, including the matcher engine (cv2.BFMatcher, FlannBasedMatcher, xfeat.XFeat, etc.), as well as the matcher_type, detector_type, descriptor_type, norm_type, and ratio_test fields. This modular structure supports extensibility and facilitates switching between traditional and learning-based feature matching backends.
The Section 5.2 reports a list of supported feature matchers.
Figure 6: Depth estimator.
Loop Detector
The diagram in Fig. 5 shows the architecture of the Loop Detector component. A central loop_detector_factory instantiates loop detectors based on the selected global_descriptor_type, which may include traditional descriptors (e.g., DBOW2, VLAD, IBOW) or deep learning-based embeddings (e.g., NetVLAD, CosPlace, EigenPlaces).
Each descriptor type creates a corresponding loop detector implementation (e.g., LoopDetectorDBoW2, LoopDetectorNetVLAD), all of which inherit from a base class hierarchy. Traditional methods inherit directly from LoopDetectorBase, while deep learning-based approaches inherit from LoopDetectorVprBase, which itself extends LoopDetectorBase. This design supports modular integration of diverse place recognition techniques within a unified loop closure framework.
The Section 5.3 reports a list of supported loop closure methods with the adopted global descriptors and local descriptor aggregation methods.
Depth Estimator
The diagram in Fig. 6 illustrates the architecture of the Depth Estimator module. A central depth_estimator_factory creates instances of various depth estimation backends based on the selected depth_estimator_type, including both traditional and learning-based methods such as DEPTH_SGBM, DEPTH_RAFT_STEREO, DEPTH_ANYTHING_V2, DEPTH_MAST3R, and DEPTH_MVDUST3R.
Each estimator type instantiates a corresponding implementation (e.g., DepthEstimatorSgbm, DepthEstimatorCrestereoMegengine, etc.), all inheriting from a common DepthEstimator interface. This base class encapsulates shared dependencies such as the camera, device, and model components, allowing for modular integration of heterogeneous depth estimation techniques across stereo, monocular, and multi-view pipelines.
The Section 5.4 reports a list of supported depth estimation/prediction models.
Figure 7: Volumetric integrator.
Volumetric Integrator
The diagram in Fig. 7 illustrates the structure of the Volumetric Integrator module. At its core, the volumetric_integrator_factory generates specific volumetric integrator instances based on the selected volumetric_integrator_type, such as TSDF and GAUSSIAN_SPLATTING.
Each type instantiates a dedicated implementation (e.g., VolumetricIntegratorTSDF, VolumetricIntegratorGaussianSplatting), which inherits from a common VolumetricIntegratorBase. This base class encapsulates key components including the camera, a keyframe_queue, and the volume, enabling flexible integration of various 3D reconstruction methods within a unified pipeline.
The Section 5.5 reports a list of supported volume integration methods.
Figure 8: Semantic Mapping.
Semantic Mapping
The diagram in Fig. 8 outlines the architecture of the Semantic Mapping module. At its core is the semantic_mapping_factory, which creates semantic mapping instances according to the selected semantic_mapping_type. Currently, the supported type is DENSE, which instantiates the SemanticMappingDense class. This class extends SemanticMappingBase and runs asynchronously in a dedicated thread to process keyframes as they become available.
SemanticMappingDense integrates semantic information into the SLAM map by leveraging per-frame predictions from a semantic segmentation model. The segmentation model is instantiated via the semantic_segmentation_factory, based on the selected semantic_segmentation_type. Supported segmentation backends include DEEPLABV3, SEGFORMER, and CLIP, each of which corresponds to a dedicated class (SemanticSegmentationDeepLabV3, SemanticSegmentationSegformer, SemanticSegmentationCLIP) inheriting from the shared SemanticSegmentationBase.
The system supports multiple semantic feature representations – such as categorical labels, probability vectors, and high-dimensional feature embeddings – and fuses them into the map using configurable methods like count-based fusion, Bayesian fusion, or feature averaging.
This modular design decouples semantic segmentation from mapping logic, enabling flexible combinations of segmentation models, datasets (e.g., NYU40, Cityscapes), and fusion strategies. It also supports customization via configuration files or programmatic APIs for dataset-specific tuning or deployment.
The Section 4.3 provides a list of supported semantic segmentation methods. A more in-depth presentation of the semantic mapping module is provided in [41].
4 Usage
Open a new terminal and start experimenting with the scripts. In each new terminal you are supposed to start with this command:
[…]
commandchars=
[…]
The file config.yaml can be used as a unique entry-point to configure the system and its global configuration parameters contained in pyslam/config_parameters.py. Further information on how to configure pySLAM are provided here.
Visual odometry
The basic Visual Odometry (VO) can be run with the following commands:
[…]
commandchars=
[…]
By default, the script processes a KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings), and its groundtruth (available in the same data/videos folder). If matplotlib windows are used, you can stop main_vo.py by clicking on one of them and pressing the key ‘Q’. As explained above, this very basic script main_vo.py strictly requires a ground truth. Now, with RGBD datasets, you can also test the RGBD odometry with the classes VisualOdometryRgbd or VisualOdometryRgbdTensor (ground truth is not required here).
Full SLAM
Similarly, you can test the full SLAM by running main_slam.py:
[…]
commandchars=
[…]
This will process the same default KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings). You can stop it by clicking on one of the opened windows and pressing the key ‘Q’ or closing the 3D pangolin GUI.
Selecting a dataset and different configuration parameters
The file config.yaml can be used as a unique entry-point to configure the system, the target dataset and its global configuration parameters set in pyslam/config_parameters.py. To process a different dataset with both VO and SLAM scripts, you need to update the file config.yaml:
  • Select your dataset type in the section DATASET (further details in the section Datasets below for further details). This identifies a corresponding dataset section (e.g. KITTI_DATASET, TUM_DATASET, etc).
  • Select the sensor_type (mono, stereo, rgbd) in the chosen dataset section.
  • Select the camera settings file in the dataset section (further details in the section Camera Settings below).
  • Set the groudtruth_file accordingly. Further details in the section Datasets below (see also the files pyslam/io/ground_truth.py and pyslam/io/convert_groundtruth_to_simple.py).
You can use the section GLOBAL_PARAMETERS of the file config.yaml to override the global configuration parameters set in pyslam/config_parameters.py. This is particularly useful when running a SLAM evaluation.
4.1 Feature tracking
If you just want to test the basic feature tracking capabilities (feature detector + feature descriptor + feature matcher) and get a taste of the different available local features, run
[…]
commandchars=
[…]
In any of the above scripts, you can choose any detector/descriptor among ORB, SIFT, SURF, BRISK, AKAZE, SuperPoint, etc. (see the section Supported Local Features below for further information).
Some basic examples are available in the subfolder test/cv. In particular, as for feature detection/description, you may want to take a look at test/cv/test_feature_manager.py too.
4.2 Loop closing
Many loop closing methods are available, combining different aggregation methods and global descriptors.
While running full SLAM, loop closing is enabled by default and can be disabled by setting kUseLoopClosing=False in pyslam/config_parameters.py. Different configuration options LoopDetectorConfigs can be found in pyslam/loop_closing/loop_detector_configs.py: Code comments provide additional useful details.
One can start experimenting with loop closing methods by using the examples in test/loopclosing. The example test/loopclosing/test_loop_detector.py is the recommended entry point.
Vocabulary management
DBoW2, DBoW3, and VLAD require pre-trained vocabularies. ORB-based vocabularies are automatically downloaded into the data folder (see pyslam/loop_closing/loop_detector_configs.py).
To create a new vocabulary, follow these steps:
  1. 1.
    Generate an array of descriptors: Use the script
    test/loopclosing/test_gen_des_array_from_imgs.py to generate the array of descriptors that will be used to train the new vocabulary. Select your desired descriptor type via the tracker configuration.
  2. 2.
    DBOW vocabulary generation: Train your target DBOW vocabulary by using the script test/loopclosing/test_gen_dbow_voc_from_des_array.py.
  3. 3.
    VLAD vocabulary generation: Train your target VLAD “vocabulary” by using the script test/loopclosing/test_gen_vlad_voc_from_des_array.py.
Once you have trained the vocabulary, you can add it in pyslam/loop_closing/loop_detector_vocabulary.py and correspondingly create a new loop detector configuration in pyslam/loop_closing/loop_detector_configs.py that uses it.
Vocabulary-free loop closing
Most methods do not require pre-trained vocabularies. Specifically:
  • iBoW and OBindex2: These methods incrementally build bags of binary words and, if needed, convert (front-end) non-binary descriptors into binary ones.
  • Others: Methods like HDC_DELF, SAD, AlexNet, NetVLAD, CosPlace, and EigenPlaces directly extract their specific global descriptors and process them using dedicated aggregators, independently from the used front-end descriptors.
As mentioned above, only DBoW2, DBoW3, and VLAD require pre-trained vocabularies.
Verify your loop detection configuration and verify vocabulary compability
Loop detection method based on a pre-trained vocabulary
When selecting a loop detection method based on a pre-trained vocabulary (such as DBoW2, DBoW3, and VLAD), ensure the following:
  1. 1.
    The back-end and the front-end are using the same descriptor type (this is also automatically checked for consistency) or their descriptor managers are independent (see further details in the configuration options LoopDetectorConfigs available in pyslam/loop_closing/loop_detector_configs.py).
  2. 2.
    A corresponding pre-trained vocubulary is available. For more details, refer to the vocabulary management section.
Missing vocabulary for the selected front-end descriptor type
If you lack a compatible vocabulary for the selected front-end descriptor type, you can follow one of these options:
1. Create and load the vocabulary (refer to the vocabulary management section).
2. Choose an *_INDEPENDENT loop detector method, which works with an independent local_feature_manager.
3. Select a vocabulary-free loop closing method.
See the file pyslam/loop_closing/loop_detector_configs.py for further details.
4.3 Volumetric reconstruction
Dense reconstruction while running SLAM
The SLAM back-end hosts a volumetric reconstruction pipeline. This is disabled by default. You can enable it by setting kUseVolumetricIntegration=True and selecting your preferred method kVolumetricIntegrationType in pyslam/config_parameters.py. At present, two methods are available: TSDF and GAUSSIAN_SPLATTING (see pyslam/dense/volumetric_integrator_factory.py). Note that you need CUDA in order to run GAUSSIAN_SPLATTING method.
At present, the volumetric reconstruction pipeline works with:
  • RGBD datasets
  • When a depth estimator is used
    • $\circ$
      in the back-end with STEREO datasets (you can’t use depth prediction in the back-end with MONOCULAR datasets, further details here)
    • $\circ$
      in the front-end (to emulate an RGBD sensor) and a depth prediction/estimation gets available for each processed keyframe.
To obtain a mesh as output, set kVolumetricIntegrationExtractMesh=True in config_parameters.py.
Reload a saved sparse map and perform dense reconstruction
Use the script main_map_dense_reconstruction.py to reload a saved sparse map and perform dense reconstruction by using its posed keyframes as input. You can select your preferred dense reconstruction method directly in the script.
  • To check what the volumetric integrator is doing, run in another shell tail -f logs/volumetric_integrator.log (from repository root folder).
  • To save the obtained dense and sparse maps, press the Save button on the GUI.
Reload and check your dense reconstruction
You can check the output pointcloud/mesh by using CloudCompare.
In the case of a saved Gaussian splatting model, you can visualize it by:
  1. 1.
    Using the superslat editor (drag and drop the saved Gaussian splatting .ply pointcloud in the editor interface).
  2. 2.
    Getting into the folder test/gaussian_splatting and running:
    \$ python test_gsm.py --load <gs_checkpoint_path>
    The directory <gs_checkpoint_path> is expected to have the following structure:
[…]
    commandchars=
[…]
Controlling the spatial distribution of keyframe FOV centers
If you are targeting volumetric reconstruction while running SLAM, you can enable a keyframe generation policy designed to manage the spatial distribution of keyframe field-of-view (FOV) centers. The FOV center of a camera is defined as the backprojection of its image center, computed using the median depth of the frame. With this policy, a new keyframe is generated only if its FOV center lies beyond a predefined distance from the nearest existing keyframe’s FOV center. You can enable this policy by setting the following parameters in the yaml setting:
[…]
commandchars=
[…]
Depth prediction
The available depth prediction models can be utilized both in the SLAM back-end and front-end.
  • Back-end: Depth prediction can be enabled in the volumetric reconstruction pipeline by setting the parameter kVolumetricIntegrationUseDepthEstimator=True and selecting your preferred kVolumetricIntegrationDepthEstimatorType in pyslam/config_parameters.py.
  • Front-end: Depth prediction can be enabled in the front-end by setting the parameter kUseDepthEstimatorInFrontEnd in pyslam/config_parameters.py. This feature estimates depth images from input color images to emulate a RGBD camera. Please, note this functionality is still experimental at present time [WIP].
Notes:
  • In the case of a monocular SLAM, do NOT use depth prediction in the back-end volumetric integration: The SLAM (fake) scale will conflict with the absolute metric scale of depth predictions. With monocular datasets, you can enable depth prediction to run in the front-end (to emulate an RGBD sensor).
  • Depth inference may be very slow (for instance, with DepthPro it takes ~1s per image on a typical machine). Therefore, the resulting volumetric reconstruction pipeline may be very slow.
Refer to the file pyslam/depth_estimation/depth_estimator_factory.py for further details. Both stereo and monocular prediction approaches are supported. You can test depth prediction/estimation by using the script main_depth_prediction.py.
Semantic mapping
The semantic mapping pipeline can be enabled by setting the parameter kDoSemanticMapping=True in config_parameters.py. The best way of configuring the semantic mapping module used is to modify it in semantic_mapping_configs.py.
Different semantic mapping methods are available Currently, we support semantic mapping using dense semantic segmenetation.
  • DEEPLABV3: from torchvision, pre-trained on COCO/VOC.
  • SEGFORMER: from transformers, pre-trained on Cityscapes or ADE20k.
  • CLIP: from f3rm package for open-vocabulary support.
Semantic features are assigned to keypoints on the image and fused into map points. The semantic features can be:
  • Labels: categorical labels as numbers.
  • Probability vectors: probability vectors for each class.
  • Feature vectors: feature vectors obtained from an encoder. This is generally used for open vocabulary mapping.
4.4 Saving and reloading
Save the a map
When you run the script main_slam.py (main_map_dense_reconstruction.py):
  • You can save the current map state by pressing the button Save on the GUI. This saves the current map along with front-end, and backend configurations into the default folder results/slam_state (results/slam_state_dense_reconstruction).
  • To change the default saving path, open config.yaml and update target folder_path in the section:
[…]
    commandchars=
[…]
Reload a saved map and relocalize in it
  • A saved map can be loaded and visualized in the GUI by running:
[…]
    commandchars=
[…]
  • To enable map reloading and relocalization when running main_slam.py, open config.yaml and set
[…]
    commandchars=
[…]
Note that pressing the Save button saves the current map, front-end, and backend configurations. Reloading a saved map replaces the current system configurations to ensure descriptor compatibility.
Trajectory saving
Estimated trajectories can be saved in three formats: TUM (The Open Mapping format), KITTI (KITTI Odometry format), and EuRoC (EuRoC MAV format). pySLAM saves two types of trajectory estimates:
  • Online: In online trajectories, each pose estimate depends only on past poses. A pose estimate is saved at the end of each front-end iteration for the current frame.
  • Final: In final trajectories, each pose estimate depends on both past and future poses. A pose estimate is refined multiple times by LBA windows that include it, as well as by PGO and GBA during loop closures.
To enable trajectory saving, open config.yaml and search for the SAVE_TRAJECTORY: set save_trajectory: True, select your format_type (tum, kitti, euroc), and the output filename. For instance for a tum format output:
[…]
commandchars=
[…]
Optimization engines
Currently, pySLAM supports both g2o and gtsam for graph optimization, with g2o set as the default engine. You can enable gtsam by setting to True the following parameters in config_parameters.py:
[…]
commandchars=
[…]
Additionally, the gtsam_factors package provides custom Python bindings for features not available in the original gtsam framework. See here for further details.
4.5 SLAM GUI
Some quick information about the non-trivial GUI buttons of main_slam.py:
  • Step: Enter the Step by step mode. Press the button Step a first time to pause. Then, press it again to make the pipeline process a single new frame.
  • Save: Save the map into the file map.json. You can visualize it back by using the script main_map_viewer.py (as explained above).
  • Reset: Reset SLAM system.
  • Draw Ground Truth: If a ground truth dataset (e.g., KITTI, TUM, EUROC, or REPLICA) is loaded, you can visualize it by pressing this button. The ground truth trajectory will be displayed in 3D and will be progressively aligned with the estimated trajectory, updating approximately every 10-30 frames. As more frames are processed, the alignment between the ground truth and estimated trajectory becomes more accurate. After about 20 frames, if the button is pressed, a window will appear showing the Cartesian alignment errors along the main axes (i.e., $e_{x},e_{y},e_{z}$ and the history of the total $RMSE$ between the ground truth and the aligned estimated trajectories.
4.6 Monitor the logs for tracking, local mapping, and loop closing simultaneously
The logs generated by the modules local_mapping.py, loop_closing.py, loop_detecting_process.py, global_bundle_adjustments.py, and volumetric integrator_<X>.py are collected in the files local_mapping.log, loop_closing.log, loop_detecting.log, gba.log, and volumetric_integrator.log, which are all stored in the folder logs.
For debugging, you can monitor one of the them in parallel by running the following command in a separate shell:
\$ tail -f logs/<log file name>
Otherwise, to check all parallel logs with tmux, run:
\$ ./scripts/launch_tmux_logs.sh
To launch slam and check all logs in a single tmux, run:
\$ ./scripts/launch_tmux_slam.sh
Press CTRL+A followed by CTRL+Q to exit from tmux environment.
4.7 Evaluating SLAM
Run a SLAM evaluation
The main_slam_evaluation.py script enables automated SLAM evaluation by executing main_slam.py across a collection of datasets and configuration presets. The main input to the script is an evaluation configuration file (e.g., evaluation/configs/evaluation.json) that specifies which datasets and presets to be used. For convenience, sample configurations for the datasets TUM, EUROC and KITTI datasets are already provided in the evaluation/configs/ directory.
For each evaluation run, results are stored in a dedicated subfolder within the results directory, containing all the computed metrics. These metrics are then processed and compared. The final output is a report, available in PDF, LaTeX, and HTML formats, that includes comparison tables summarizing the Absolute Trajectory Error (ATE), the maximum deviation from the ground truth trajectory and other metrics.
Evaluation results are presented in the following subsections. For convenience, they are also available at this link. The evaluation uses commit f7c9a13 with the following presets: baseline, root_sift, and superpoint, which correspond to local feature extractors ORB2, root_SIFT, and SuperPoint, respectively.
TUM Evaluation
Table 1: TUM: Table RMSE
Dataset baseline root_sift superpoint
rgbd_dataset_freiburg1_desk 0.06634 0.05289 0.04815
rgbd_dataset_freiburg1_desk2 0.04645 0.06255 0.03868
rgbd_dataset_freiburg1_room 0.05885 0.07968 0.07199
rgbd_dataset_freiburg1_xyz 0.01196 0.01758 0.01726
rgbd_dataset_freiburg3_long_office_household 0.00921 0.00982 0.00952
rgbd_dataset_freiburg3_nostructure_texture_far 0.08397 0.11342 0.12118
rgbd_dataset_freiburg3_nostructure_texture_near_withloop 0.02245 0.05022 0.05037
Average 0.04275 0.05517 0.05102
Std Dev 0.0362 0.04531 0.04503
Best (Average) Preset baseline
Best (Average) Metric 0.04275
Table 2: TUM: Table Max
Dataset baseline root_sift superpoint
rgbd_dataset_freiburg1_desk 0.326 0.19361 0.22686
rgbd_dataset_freiburg1_desk2 0.44543 0.28865 0.38157
rgbd_dataset_freiburg1_room 0.14959 0.2838 0.17632
rgbd_dataset_freiburg1_xyz 0.06059 0.0677 0.0635
rgbd_dataset_freiburg3_long_office_household 0.03609 0.03543 0.04005
rgbd_dataset_freiburg3_nostructure_texture_far 0.29727 0.4361 0.70826
rgbd_dataset_freiburg3_nostructure_texture_near_withloop 0.09011 0.13774 0.11602
Average 0.20073 0.20615 0.24465
Std Dev 0.18661 0.18182 0.26334
Best (Average) Preset baseline
Best (Average) Metric 0.20073
Table 3: TUM: Table Percent Lost
Dataset baseline root_sift superpoint
rgbd_dataset_freiburg1_desk 1.364 1.502 1.674
rgbd_dataset_freiburg1_desk2 0.548 1.55 0.612
rgbd_dataset_freiburg1_room 0.074 0.148 0.088
rgbd_dataset_freiburg1_xyz 0.05 0.076 0.078
rgbd_dataset_freiburg3_long_office_household 0.0 0.0 0.0
rgbd_dataset_freiburg3_nostructure_texture_far 0.622 0.622 0.666
rgbd_dataset_freiburg3_nostructure_texture_near_withloop 0.024 0.036 0.036
Average 0.38314 0.562 0.45057
Std Dev 0.50438 0.82947 0.59847
Best (Average) Preset baseline
Best (Average) Metric 0.38314
EUROC Evaluation
Table 4: EUROC: Table RMSE
Dataset baseline root_sift superpoint
V101 0.08819 0.08816 0.08834
V102 0.06852 0.06609 0.06694
V201 0.07062 0.06409 0.09735
V202 0.62539 0.56925 0.53736
MH01 0.03775 0.03781 0.04172
MH02 0.04217 0.0456 0.04177
MH03 0.0512 0.04905 0.04882
MH04 0.0702 0.05907 0.0646
MH05 0.08177 0.06607 0.07131
Average 0.1262 0.11613 0.11758
Std Dev 0.19831 0.17517 0.15435
Best (Average) Preset root_sift
Best (Average) Metric 0.11613
Table 5: EUROC: Table Max
Dataset baseline root_sift superpoint
V101 0.16335 0.163 0.16208
V102 0.14545 0.12379 0.13136
V201 0.15864 0.18921 0.22163
V202 2.06607 2.52628 1.97847
MH01 0.09141 0.10259 0.21303
MH02 0.11922 0.1318 0.11196
MH03 0.19995 0.17533 0.14082
MH04 0.32021 0.27938 0.26722
MH05 0.24828 0.17474 0.18991
Average 0.39029 0.42957 0.37961
Std Dev 0.67128 0.75543 0.6093
Best (Average) Preset superpoint
Best (Average) Metric 0.37961
Table 6: EUROC: Table Percent Lost
Dataset baseline root_sift superpoint
V101 0.0 0.0 0.0
V102 0.0 0.0 0.0
V201 0.04 0.04 0.04
V202 0.146 0.138 0.112
MH01 0.05 0.074 0.08
MH02 0.018 0.026 0.012
MH03 0.046 0.03 0.032
MH04 0.0 0.01 0.0
MH05 0.07 0.016 0.034
Average 0.04111 0.03711 0.03444
Std Dev 0.05351 0.04829 0.04435
Best (Average) Preset superpoint
Best (Average) Metric 0.03444
KITTI Evaluation
Table 7: KITTI: Table RMSE
Dataset baseline root_sift superpoint
00 3.947 3.24849 3.93432
01 21.60014 21.86468 23.37292
02 6.1194 5.6707 10.14455
03 5.96451 5.92601 5.88895
04 1.74556 1.76121 1.65506
05 2.13915 2.11047 2.52263
06 2.74011 2.73981 2.6784
07 1.12358 1.12823 1.09496
08 3.99078 4.13702 4.14543
09 4.26068 4.35988 4.06244
10 2.22097 2.25127 2.27694
Average 5.07744 5.01798 5.61605
Std Dev 5.47265 5.63609 6.77
Best (Average) Preset root_sift
Best (Average) Metric 5.01798
Table 8: KITTI: Table Max
Dataset baseline root_sift superpoint
00 6.82983 5.8663 7.04067
01 33.32903 33.85867 36.62253
02 13.97623 9.47974 21.05188
03 10.16642 10.16978 10.06014
04 3.09135 3.0576 2.87271
05 3.81484 3.77687 4.9547
06 6.53758 5.14239 5.11673
07 1.84418 1.82355 1.8655
08 10.809 11.23635 10.75959
09 8.53888 8.70237 8.31166
10 4.13901 4.37849 4.27905
Average 9.37058 8.86292 10.26683
Std Dev 8.62548 8.62145 12.23529
Best (Average) Preset root_sift
Best (Average) Metric 8.86292
Table 9: KITTI: Table Percent Lost
Dataset baseline root_sift superpoint
00 0.0 0.038 0.012
01 0.0 0.0 0.0
02 0.012 0.004 0.004
03 0.0 0.0 0.0
04 0.0 0.0 0.0
05 0.0 0.0 0.028
06 0.108 0.018 0.0
07 0.0 0.0 0.0
08 0.0 0.0 0.0
09 0.0 0.0 0.0
10 0.0 0.0 0.0
Average 0.01091 0.00545 0.004
Std Dev 0.07244 0.02373 0.0196
Best (Average) Preset superpoint
Best (Average) Metric 0.004
pySLAM performances and comparative evaluations
For a comparative evaluation of the “online” trajectory estimated by pySLAM versus the “final” trajectory estimated by ORB-SLAM3, check out this nice notebook. For more details about “online” and “final” trajectories, refer to Section 4.4.
Note: Unlike ORB-SLAM3, which only saves the final pose estimates (recorded after the entire dataset has been processed), pySLAM saves both online and final pose estimates. For details on how to save trajectories in pySLAM, refer to this section. When you click the Draw Ground Truth button in the GUI (see here), you can visualize the Absolute Trajectory Error (ATE or RMSE) history and evaluate both online and final errors up to the current time.
5 Supported components and models
5.1 Supported local features
Figure 9: Timeline of some of the most famous local features for image matching, place recognition and SLAM.
Fig. 9 shows a timeline with some of the most famous local features for image matching, place recognition and SLAM. At present time, pySLAM supports the following feature detectors:
The following feature descriptors are supported:
For more information, refer to pyslam/local_features/feature_types.py file. Some of the local features consist of a joint detector-descriptor. You can start playing with the supported local features by taking a look at test/cv/test_feature_manager.py and main_feature_matching.py.
In both the scripts main_vo.py and main_slam.py, you can create your preferred detector-descritor configuration and feed it to the function feature_tracker_factory(). Some ready-to-use configurations are already available in the file pyslam/local_features/feature_tracker.configs.py
The function feature_tracker_factory() can be found in the file pyslam/local_features/feature_tracker.py. Take a look at the file pyslam/local_features/feature_manager.py for further details.
N.B.: You just need a single python environment to be able to work with all the supported local features!
5.2 Supported matchers
See the file pyslam/local_features/feature_matcher.py for further details.
5.3 Supported global descriptors and local descriptor aggregation methods
Local descriptor aggregation methods
NOTE: iBoW and OBIndex2 incrementally build a binary image index and do not need a prebuilt vocabulary. In the implemented classes, when needed, the input non-binary local descriptors are transparently transformed into binary descriptors.
Global descriptors
Also referred to as holistic descriptors:
Different loop closing methods are available. These combines the above aggregation methods and global descriptors. See the file pyslam/loop_closing/loop_detector_configs.py for further details.
5.4 Supported depth prediction models
Both monocular and stereo depth prediction models are available. SGBM algorithm has been included as a classic reference approach.
5.5 Supported volumetric mapping methods
  • TSDF with voxel block grid (parallel spatial hashing) [15]
  • Incremental 3D Gaussian Splatting. See here and MonoGS for a description of its backend [37, 21].
Supported semantic segmentation methods
  • DeepLabv3 [13]: from torchvision, pre-trained on COCO/VOC.
  • Segformer [70]: from transformers, pre-trained on Cityscapes or ADE20k.
  • CLIP [28]: from f3rm package for open-vocabulary support.
5.6 Configuration
Main configuration file
Refer to this section for how to update the main configuration file config.yaml and affect the configuration parameters in config_parameters.py.
Datasets
The following datasets are supported:
Dataset type in config.yaml
KITTI odometry data set (grayscale, 22 GB) type: KITTI_DATASET
TUM dataset type: TUM_DATASET
ICL-NUIM dataset type: ICL_NUIM_DATASET
EUROC dataset type: EUROC_DATASET
REPLICA dataset type: REPLICA_DATASET
TARTANAIR dataset type: TARTANAIR_DATASET
ScanNet dataset type: SCANNET_DATASET
ROS1 bags type: ROS1BAG_DATASET
ROS2 bags type: ROS2BAG_DATASET
Video file type: VIDEO_DATASET
Folder of images type: FOLDER_DATASET
Use the download scripts available in the folder scripts to download some of the following datasets.
KITTI Datasets
pySLAM code expects the following structure in the specified KITTI path folder (specified in the section KITTI_DATASET of the file config.yaml). :
[…]
commandchars=
[…]
  1. 1.
    Download the dataset (grayscale images) from http://www.cvlibs.net/datasets/kitti/eval_odometry.php and prepare the KITTI folder as specified above
  2. 2.
    Select the corresponding calibration settings file (section KITTI_DATASET: settings: in the file config.yaml)
TUM Datasets
pySLAM code expects a file associations.txt in each TUM dataset folder (specified in the section TUM_DATASET: of the file config.yaml).
  1. 1.
    Download a sequence from vision.in.tum.de/data/datasets/rgbd-dataset/download and uncompress it.
  2. 2.
    Associate RGB images and depth images using the python script associate.py. You can generate your associations.txt file by executing:
    \$ python associate.py PATH_TO_SEQUENCE/rgb.txt PATH_TO_SEQUENCE/depth.txt > associations.txt      # pay attention to the order!
  3. 3.
    Select the corresponding calibration settings file (section TUM_DATASET: settings: in the file config.yaml).
ICL-NUIM Datasets
Follow the same instructions provided for the TUM datasets.
EuRoC Datasets
  1. 1.
    Download a sequence (ASL format) from http://projects.asl.ethz.ch/datasets/doku.php?id=kmavvisualinertialdatasets (check this direct link)
  2. 2.
    Use the script io/generate_euroc_groundtruths_as_tum.sh to generate the TUM-like groundtruth files path + '/' + name + '/mav0/state_groundtruth_estimate0/data.tum' that are required by the EurocGroundTruth class.
  3. 3.
    Select the corresponding calibration settings file (section EUROC_DATASET: settings: in the file config.yaml).
Replica Datasets
  1. 1.
    You can download the zip file containing all the sequences by running:
    \$ wget https://cvg-data.inf.ethz.ch/nice-slam/data/Replica.zip
  2. 2.
    Then, uncompress it and deploy the files as you wish.
  3. 3.
    Select the corresponding calibration settings file (section REPLICA_DATASET: settings: in the file config.yaml).
Tartanair Datasets
  1. 1.
    You can download the datasets from https://theairlab.org/tartanair-dataset/
  2. 2.
    Then, uncompress them and deploy the files as you wish.
  3. 3.
    Select the corresponding calibration settings file (section TARTANAIR_DATASET: settings: in the file config.yaml).
ScanNet Datasets
  1. 1.
    You can download the datasets following instructions in http://www.scan-net.org. You will need to request the dataset from the authors.
  2. 2.
    There are two versions you can download:
    • A subset of pre-processed data termed as tasks/scannet_frames_2k: this version is smaller, and more generally available for training neural networks. However, it only includes one frame out of each 100, which makes it unusable for SLAM. The labels are processed by mapping them from the original Scannet label annotations to NYU40.
    • The raw data: this version is the one used for SLAM. You can download the whole dataset (TBs of data) or specific scenes. A common approach for evaluation of semantic mapping is to use the scannetv2_val.txt scenes. For downloading and processing the data, you can use the following repository as the original Scannet repository is tested under Python 2.7 and does’t support batch downloading of scenes.
  3. 3.
    Once you have the color, depth, pose, and (optional for semantic mapping) label folders, you should place them following {path_to_scannet}/scans/{scene_name}/[color, depth, pose, label]. Then, configure the base_path and name in the file config.yaml.
  4. 4.
    Select the corresponding calibration settings file (section SCANNET_DATASET: settings: in the file config.yaml). NOTE: the RGB images are rescaled to match the depth image. The current intrinsic parametes in the existing calibration file reflect that.
ROS1 bags
  1. 1.
    Source the main ROS1 setup.bash after you have sourced the pyslam python environment.
  2. 2.
    Set the paths and ROS1BAG_DATASET: ros_parameters in the file config.yaml.
  3. 3.
    Select/prepare the correspoding calibration settings file (section ROS1BAG_DATASET: settings: in the file config.yaml). See the available yaml files in the folder Settings as an example.
ROS2 bags
  1. 1.
    Source the main ROS2 setup.bash after you have sourced the pyslam python environment.
  2. 2.
    Set the paths and ROS2BAG_DATASET: ros_parameters in the file config.yaml.
  3. 3.
    Select/prepare the correspoding calibration settings file (section ROS2BAG_DATASET: settings: in the file config.yaml). See the available yaml files in the folder Settings as an example.
Video and Folder Datasets
You can use the VIDEO_DATASET and FOLDER_DATASET types to read generic video files and image folders (specifying a glob pattern), respectively. A companion ground truth file can be set in the simple format type: Refer to the class SimpleGroundTruth in io/ground_truth.py and check the script io/convert_groundtruth_to_simple.py.
5.7 Camera Settings
The folder settings contains the camera settings files which can be used for testing the code. These are the same used in the framework ORB-SLAM2 [43]. You can easily modify one of those files for creating your own new calibration file (for your new datasets).
In order to calibrate your camera, you can use the scripts in the folder calibration. In particular: 1. Use the script grab_chessboard_images.py to collect a sequence of images where the chessboard can be detected (set the chessboard size therein, you can use the calibration pattern calib_pattern.pdf in the same folder) 2. Use the script calibrate.py to process the collected images and compute the calibration parameters (set the chessboard size therein)
For more information on the calibration process, see this tutorial [35] or this other link [49].
If you want to use your camera, you have to:
  • Calibrate it and configure settings/WEBCAM.yaml accordingly.
  • Record a video (for instance, by using save_video.py in the folder calibration).
  • Configure the VIDEO_DATASET section of config.yaml in order to point to your recorded video.
6 Credits
The following is a list of frameworks that inspired or has been integrated into pySLAM. Many thanks to their Authors for their great work.